# Improved quantum data analysis

### 12 Dec 2023 v3

@@ 2.4 Naive expectation estimation
$f(t)$ =$=\begin{cases}1,&\lvert t-\tau\rvert\geq c,\\ 0,&\text{otherwise}.\end{cases}$
@@
If $E$ is a projector, then $A_{k}$ is a projector and $A_{k}A_{\ell}=A_{\ell}A_{k}=0$ for all $k,\ell\in\{0,\dotsc,n\}$. Since $B$ is a sum of orthogonal projectors $A_{k}$ with $k\in\{0,\dotsc,n\}$, it follows that $B$ is a projector. ∎
@@ 2.5 Quantum union bound-style results
The following result is part of the “Damage Lemma” of Aaronson and Rothblum [AR19, Lemma 17]. Since the original proof of the “Damage Lemma” was found to be incorrect [Lei22], we provide a slightly different proof by induction below:
@@
$\lvert\operatorname{tr}(S_{m}(\dotsb S_{1}(\rho)))-p_{[m]}\rvert$ $\leq 2\cdot\sum_{k=1}^{m-1}p_{[k]}\cdot d_{\mathrm{tr}}(\sigma_{k},\rho).\qed$
@@
The following inequality, which appears in the proof of [OV22, Theorem 1.3], will be used to show that when $S_{1},\dotsc,S_{m}$ are applied sequentially to the initial state $\rho$, the probability of observing $S_{1},\dotsc,S_{t-1}$ accept and $S_{t}$ reject for certain “good” values of $t\in[m]$ is bounded below by a positive constant for specific $\rho$ and $S_{1},\dotsc,S_{m}$ (see proof of Lemma 4.2).
@@
By [OV22, Lemma 2.1] and the inequality $\operatorname{\mathbbm{1}}-\sqrt{A_{i}}\leq\overline{A_{i}}$,
@@
Finally, for the “unique decoding” part of our Hypothesis Selection routine we will use a related result, Gao’s quantum Union Bound [Gao15]:
@@ 3 $\chi^{2}$-stable Threshold Reporting
The following is well known [Ber95]:
@@
Proof.
Let ${\boldsymbol{S}}^{\prime}$ be an independent copy of ${\boldsymbol{S}}$. Since the function $f$ is 1-Lipschitz, we always have $\tfrac{1}{2}(f({\boldsymbol{S}})-f({\boldsymbol{S}}^{\prime}))^{2}\leq\tfrac{1}{2}({\boldsymbol{S}}-{\boldsymbol{S}}^{\prime})^{2}$. The result follows by taking expectations of both sides. ∎
@@
But this indeed holds with $C=(e-1)^{2}$, as it is equivalent to $e^{\lambda}\leq 1+(e-1)\lambda$ on $[0,1]$, which follows from convexity of $\lambda\mapsto e^{\lambda}$. ∎
@@
where the last step uses $\operatorname*{\mathbf{E}}[f({\boldsymbol{S}})]=\operatorname*{\mathbf{P}}[B]$. ∎
@@
Theorem (Theorem 1.2, restated).
Fix a threshold $\theta\in[0,1]$. Let ${\boldsymbol{S}}\sim\operatorname{Binomial}(n,p)$ and write $q=1-p$. Assume that ${\boldsymbol{X}}$ is an independent Exponential random variable with mean at least $\operatorname*{\mathbf{stddev}}[{\boldsymbol{S}}]=\sqrt{pqn}$ (and also at least $1$). Let $B$ be the event that ${\boldsymbol{S}}+{\boldsymbol{X}}>\theta n$, and assume that $\operatorname*{\mathbf{P}}[B]<\frac{1}{4}$. Then
$d_{\mathrm{\chi^{2}}}(({\boldsymbol{S}}\mid\overline{B}),{\boldsymbol{S}})\lesssim\left\lparen\operatorname*{\mathbf{P}}[B]\cdot\frac{\operatorname*{\mathbf{stddev}}[{\boldsymbol{S}}]}{\operatorname*{\mathbf{E}}[{\boldsymbol{X}}]}\right\rparen^{2}.$
Proof.
@@
Intuitively this holds because $g(s)$ should not be much different than $f(s)$, and $\operatorname*{\mathbf{E}}[f({\boldsymbol{S}})]=\operatorname*{\mathbf{P}}[B]$ by definition. Formally, we consider two cases: $p\geq\frac{1}{n}$ (intuitively, the main case) and $p\leq\frac{1}{n}$.
@@
$p\geq\frac{1}{n}$. In this case we use that $\operatorname*{\mathbf{P}}[{\boldsymbol{S}}>pn]\geq\frac{1}{4}$ (see, e.g., [Doe18]), and hence: (i) it must be that $\theta\geq p$, since we are assuming $\operatorname*{\mathbf{P}}[B]=\operatorname*{\mathbf{P}}[{\boldsymbol{S}}+{\boldsymbol{X}}>\theta n]<\frac{1}{4}$; and, (ii) $\operatorname*{\mathbf{P}}[B]\geq\operatorname*{\mathbf{P}}[{\boldsymbol{S}}>pn]\cdot\operatorname*{\mathbf{P}}[{\boldsymbol{X}}\geq(\theta-p)n]\geq\frac{1}{4}\exp(-\lambda(\theta-p)n)$, where the first inequality used independence of ${\boldsymbol{S}}$ and ${\boldsymbol{X}}$ and the second inequality used $(\theta-p)n\geq 0$ (by (i)). Thus to establish Equation 3, it remains to show $\exp(-\lambda\theta n)\cdot(q+pe^{\lambda})^{n}\lesssim\exp(-\lambda(\theta-p)n)$.
@@
$(q+pe^{\lambda})^{n}$ $=(1+p(e^{\lambda}-1))^{n}\leq\exp(p(e^{\lambda}-1)n)\leq\exp(\lambda pn)\exp(e\lambda^{2}\cdot p\cdot n)\qquad\text{and}$
@@
$p\leq\frac{1}{n}$. Since $\lambda\in(0,1]$, we have $e^{\lambda}\leq 1+2\lambda$. Hence, $q+pe^{\lambda}\leq 1+2p\lambda\leq 1+\frac{2}{n}$, and so $(q+pe^{\lambda})^{n}\lesssim 1$, meaning that Equation 3 follows from $\operatorname*{\mathbf{P}}[B]\geq\operatorname*{\mathbf{P}}[{\boldsymbol{X}}>\theta n]=\exp(-\lambda\theta n)$. ∎
@@ 3.1 The quantum version
The statistics of the measurement $\{E_{k}\mid k=0,\dotsc,n\}$ applied to $\varrho$ are distributed as $\operatorname{Binomial}(n,\operatorname{tr}(\rho A))$, so $\operatorname*{\mathbf{E}}_{\varrho}[E_{k}]=\operatorname*{\mathbf{P}}[{\boldsymbol{S}}=k]$. Hence,
@@
$\operatorname{F}\left\lparen\varrho,{\left.\kern-1.2pt{\varrho}\right|}_{\sqrt{{\mathbbm{1}}-B}}\right\rparen$ ) $=\sum_{k=0}^{n}\sqrt{\operatorname{tr}(\varrho\cdot E_{k})}\sqrt{\operatorname{tr}\left\lparen{\left.\kern-1.2pt{\varrho}\right|}_{\sqrt{{\mathbbm{1}}-B}}\cdot E_{k}\right\rparen}=\operatorname{BC}(({\boldsymbol{S}}\mid{\boldsymbol{S}}+{\boldsymbol{X}}\leq\theta n),{\boldsymbol{S}}).\qed$
@@ 4.1 Preliminary reductions
Let $\rho\in\mathbb{C}^{d\times d}$ denote the unknown quantum state and let $A_{1},\dotsc,A_{m}\in\mathbb{C}^{d\times d}$ be the observables in the quantum Threshold Search problem (which we assume are given in an online fashion). If we extend the unknown state $\rho$ to $\rho\otimes\lvert 0\rangle\langle 0\rvert$, then by Naimark’s Theorem 2.1, there exists a projector $\Pi_{i}\in\mathbb{C}^{d\times d}\otimes\mathbb{C}^{2\times 2}$ for each $A_{i}$ such that $\operatorname*{\mathbf{E}}_{\rho\otimes\lvert 0\rangle\langle 0\rvert}[\Pi_{i}]=\operatorname*{\mathbf{E}}_{\rho}[A_{i}]$ for all $i=1,\dotsc,m$. Since the state $\rho\otimes\lvert 0\rangle\langle 0\rvert$ can be prepared without knowing $\rho$ and this extension increases the dimension of the quantum system only by a constant factor, by replacing $\rho$ by $\rho\otimes\lvert 0\rangle\langle 0\rvert$ and each $A_{i}$ by the corresponding $\Pi_{i}$, it follows that we can assume, without loss of generality, that the observables $A_{1},\dotsc,A_{m}$ are projectors.
@@
Proof of Theorem 4.1, assuming Lemma 4.2..
We will use the algorithm in Lemma 4.2 as a kind of “subroutine” for the main theorem. Our first step is to augment this subroutine in the following way:
@@
We want to show that the algorithm does not make errors of type FP or FN with probability at least $0.1$. To this end, we introduce the following notation.
Notation 4.3.
For $i=1,\dotsc,m$, let:
@@
Therefore, since $53/4<\exp(-1/4)<4/$ 5, we have $q_{t}\leq 4/5$ and $q_{t-1}\geq 3/4$, for $\operatorname*{\mathbf{E}}[{\boldsymbol{X}}]=\Omega(\sqrt{n})$. ∎
Assuming $\operatorname*{\mathbf{E}}[{\boldsymbol{X}}]=\Omega(\sqrt{n})$, let $t\in[m]$ be as in 4.4. Since $q_{m}\leq q_{t}\leq 4/5$, it follows that the probability the algorithm makes an FN error is at most $4/5$. In fact, since $q_{t}\leq 4/5$, the algorithm will pick an index $i\leq t$ with probability at least $1/5$. Thus, to show that the algorithm succeeds w.p. at least $0.1$, it suffices to show that w.h.p. the algorithm does not pick an index $i\in\mathcal{B}$, where $\mathcal{B}\subseteq[m]$ is the subset defined by
@@
Proof of Lemma 4.2.
Fix $\eta=0.01$, so that indeed $n=O(\log^{2}m)$ as promised. By Lemma 2.6,
@@
By 4.5 and the Cauchy–Schwarz inequality,
@@
where $i\not\in\mathcal{B}$ denotes $i\in[t]\setminus\mathcal{B}$. By 4.4, $p_{1}+\dotsb+p_{t}\leq 1/4$. Hence,
@@
Since $\operatorname{F}(\rho^{\otimes n},\varrho_{t})\leq 1$, $\eta=0.01$, and, by 4.4, $q_{t}\leq 4/5$, it follows that
@@
Since $\sum_{i\not\in\mathcal{B}}s_{i}$ is the probability that the algorithm returns an index $i\in[t]$ with $\operatorname*{\mathbf{E}}_{\rho}[A_{i}]\geq 1/3$, it follows that the algorithm is correct with probability at least $0.03$. ∎
@@ 5.1 Shadow Tomography
We begin by describing how to deduce our online Shadow Tomography result, Theorem 1.4, from our online Threshold Search result, Theorem 1.1. As mentioned earlier, this deduction is known [Aar20] to follow almost immediately from a mistake-bounded learning algorithm for quantum states due to Aaronson, Chen, Hazan, Kale, and Nayak [${}^{+}$19], described below. We will fill in a few details that are not spelled out in [Aar20].
@@
The above theorem is similar to, but not quite the same, as “Theorem 1” in [${}^{+}$19]. However it is easy to check that [${}^{+}$19]’s Section 3.3 (“Proof of Theorem 1”) applies equally well to establish Theorem 5.1 above. [^5] Briefly: the RTFL/MMW algorithm will only do an update in the “mistake” rounds. The loss is taken to be $|\widehat{\mu}_{t}-\mu^{\prime}_{t}|$. On any mistake, we have $|\widehat{\mu}_{t}-\mu_{t}|>\frac{3}{4}\epsilon$ and $|\mu^{\prime}_{t}-\mu_{t}|\leq\frac{1}{4}\epsilon$, hence the student incurs loss at least $\frac{1}{2}\epsilon$. On the other hand, answering according to the true $\mu_{t}$ would only incur loss at most $\frac{1}{4}\epsilon$. The regret calculation bounding the number of mistakes is now the same.
@@
$\operatorname*{\mathbf{E}}_{\rho}[A_{i}]\leq\theta_{i}+\epsilon\qquad\text{and}\qquad\operatorname*{\mathbf{E}}_{\rho}[{\mathbbm{1}}-A_{i}]\leq 1-\theta_{i}+\epsilon$
@@
for some particular $j$. Then our algorithm can correctly output “ $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{j}]-\theta_{j}\rvert>\frac{3}{4}\epsilon$”. Furthermore, at this point the algorithm may use the holdout copies of $\rho$ to obtain an estimate $\mu^{\prime}_{j}$ of $\operatorname*{\mathbf{E}}_{\rho}[A_{j}]$ (in the naive way) that satisfies $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{j}]-\mu^{\prime}_{j}\rvert\leq\frac{1}{4}\epsilon$ except with probability at most $\delta/2$, provided $c$ is large enough. ∎
@@
To begin, recall that our Shadow Tomography algorithm receives quantum events $A_{1},A_{2},\dots$ in an online fashion. As it receives them, it will run the following online algorithms concurrently:
@@
The algorithm simulates both the teacher and student roles of the mistake-bounded setting of [${}^{+}$19] and runs in rounds. A new round is started whenever the teacher declares a mistake and a fresh batch of $n_{0}$ copies of the state $\rho$ is used by the teacher in each round. When it receives input $A_{t}$, the algorithm runs the next iteration of the mistake-bounded learning algorithm of Theorem 5.1 to get the student’s prediction $\widehat{\mu}_{t}$. Then it runs the next iteration of the Threshold Search algorithm from Lemma 5.2 with input ($A_{t}$, $\widehat{\mu}_{t}$); the estimates $\widehat{\mu}_{t}$ output by the student serve as the $\theta_{t}$ threshold values used in Lemma 5.2.
@@
It remains to verify that — assuming correct answers from all uses of Lemma 5.2 — our Shadow Tomography algorithm satisfies the Teacher Properties of Theorem 5.1 and also that all $m$ estimates for $\operatorname*{\mathbf{E}}_{\rho}[A_{i}]$ produced by the algorithm are correct to within an additive error $\epsilon$. Let us first note that within each round of the Shadow Tomography algorithm, we never supply more than $m$ quantum events to the Threshold Search algorithm from Lemma 5.2. The main point to observe is that if our Threshold Search routine from Lemma 5.2 ever passes on some $(A_{t},\widehat{\mu}_{t})$ pair, it must be that $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{t}]-\widehat{\mu}_{t}\rvert\leq\epsilon$; the reason is that passing implies the Threshold Search algorithm is prepared to output “ $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{i}]-\theta_{i}\rvert\leq\epsilon$ for all $i$”. On the other hand, it’s immediate from Lemma 5.2 that if the teacher declares “mistake” on some $(A_{t},\widehat{\mu}_{t})$ pair, then indeed we have $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{t}]-\widehat{\mu}_{t}\rvert>\frac{3}{4}\epsilon$, and the supplied correction $\mu^{\prime}_{t}$ satisfies $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{t}]-\mu^{\prime}_{t}\rvert\leq\frac{1}{4}\epsilon$ (as is necessary for the Teacher Properties, and is more than sufficient for the Shadow Tomography guarantee). ∎

### 15 Mar 2024 v4

@@ 2.4 Naive expectation estimation
[…]
@@
If $E$ is a projector, then $A_{k}$ is a projector and $A_{k}A_{\ell}=A_{\ell}A_{k}=0$ for all $k,\ell\in\{0,\dotsc,n\}$. Since $B$ is a sum of orthogonal projectors $A_{k}$ with $k\in\{0,\dotsc,n\}$, it follows that $B$ is a projector.
@@ 2.5 Quantum union bound-style results
The following result is part of the “Damage Lemma” of Aaronson and Rothblum [4, Lemma 17]. Since the original proof of the “Damage Lemma” was found to be incorrect [29], we provide a slightly different proof by induction below:
@@
$\lvert\operatorname{tr}(S_{m}(\dotsb S_{1}(\rho)))-p_{[m]}\rvert$ $\leq 2\cdot\sum_{k=1}^{m-1}p_{[k]}\cdot d_{\mathrm{tr}}(\sigma_{k},\rho).$
@@
The following inequality, which appears in the proof of [33, Theorem 1.3], will be used to show that when $S_{1},\dotsc,S_{m}$ are applied sequentially to the initial state $\rho$, the probability of observing $S_{1},\dotsc,S_{t-1}$ accept and $S_{t}$ reject for certain “good” values of $t\in[m]$ is bounded below by a positive constant for specific $\rho$ and $S_{1},\dotsc,S_{m}$ (see proof of Lemma 4.2).
@@
By [33, Lemma 2.1] and the inequality $\operatorname{\mathbbm{1}}-\sqrt{A_{i}}\leq\overline{A_{i}}$,
@@
Finally, for the “unique decoding” part of our Hypothesis Selection routine we will use a related result, Gao’s quantum Union Bound [20]:
@@ 3 $\chi^{2}$-stable Threshold Reporting
The following is well known [8]:
@@
Proof 3.2.
Let ${\boldsymbol{S}}^{\prime}$ be an independent copy of ${\boldsymbol{S}}$. Since the function $f$ is 1-Lipschitz, we always have $\tfrac{1}{2}(f({\boldsymbol{S}})-f({\boldsymbol{S}}^{\prime}))^{2}\leq\tfrac{1}{2}({\boldsymbol{S}}-{\boldsymbol{S}}^{\prime})^{2}$. The result follows by taking expectations of both sides.
@@
But this indeed holds with $C=(e-1)^{2}$, as it is equivalent to $e^{\lambda}\leq 1+(e-1)\lambda$ on $[0,1]$, which follows from convexity of $\lambda\mapsto e^{\lambda}$.
@@
where the last step uses $\operatorname*{\mathbf{E}}[f({\boldsymbol{S}})]=\operatorname*{\mathbf{P}}[B]$.
@@
See 1.2
Proof 3.7.
@@
Intuitively this holds because $g(s)$ should not be much different from $f(s)$, and $\operatorname*{\mathbf{E}}[f({\boldsymbol{S}})]=\operatorname*{\mathbf{P}}[B]$ by definition. Formally, we consider two cases: $p\geq\frac{1}{n}$ (intuitively, the main case) and $p\leq\frac{1}{n}$.
@@
$p\geq\frac{1}{n}$. In this case we use that $\operatorname*{\mathbf{P}}[{\boldsymbol{S}}>pn]\geq\frac{1}{4}$ (see, e.g., [16]), and hence: (i) it must be that $\theta\geq p$, since we are assuming $\operatorname*{\mathbf{P}}[B]=\operatorname*{\mathbf{P}}[{\boldsymbol{S}}+{\boldsymbol{X}}>\theta n]<\frac{1}{4}$; and, (ii) $\operatorname*{\mathbf{P}}[B]\geq\operatorname*{\mathbf{P}}[{\boldsymbol{S}}>pn]\cdot\operatorname*{\mathbf{P}}[{\boldsymbol{X}}\geq(\theta-p)n]\geq\frac{1}{4}\exp(-\lambda(\theta-p)n)$, where the first inequality used independence of ${\boldsymbol{S}}$ and ${\boldsymbol{X}}$ and the second inequality used $(\theta-p)n\geq 0$ (by (i)). Thus, to establish Equation 3, it remains to show $\exp(-\lambda\theta n)\cdot(q+pe^{\lambda})^{n}\lesssim\exp(-\lambda(\theta-p)n)$.
@@
[…]
@@
$p\leq\frac{1}{n}$. Since $\lambda\in(0,1]$, we have $e^{\lambda}\leq 1+2\lambda$. Hence, $q+pe^{\lambda}\leq 1+2p\lambda\leq 1+\frac{2}{n}$, and so $(q+pe^{\lambda})^{n}\lesssim 1$, meaning that Equation 3 follows from $\operatorname*{\mathbf{P}}[B]\geq\operatorname*{\mathbf{P}}[{\boldsymbol{X}}>\theta n]=\exp(-\lambda\theta n)$.
@@ 3.1 The quantum version
The statistics of the measurement $\{E_{k}\mid k=0,\dotsc,n\}$ applied to $\varrho$ follow a binomial distribution $\operatorname{Binomial}(n,\operatorname{tr}(\rho A))$, so $\operatorname*{\mathbf{E}}_{\varrho}[E_{k}]=\operatorname*{\mathbf{P}}[{\boldsymbol{S}}=k]$. Hence,
@@
$\operatorname{F}\left\lparen\varrho,{\left.\kern-1.2pt{\varrho}\right|}_{\sqrt{{\mathbbm{1}}-B}}\right\rparen$ ) $=\sum_{k=0}^{n}\sqrt{\operatorname{tr}(\varrho\cdot E_{k})}\sqrt{\operatorname{tr}\left\lparen{\left.\kern-1.2pt{\varrho}\right|}_{\sqrt{{\mathbbm{1}}-B}}\cdot E_{k}\right\rparen}=\operatorname{BC}(({\boldsymbol{S}}\mid{\boldsymbol{S}}+{\boldsymbol{X}}\leq\theta n),{\boldsymbol{S}}).$ ) .
@@ 4.1 Preliminary reductions
Let $\rho\in\mathbb{C}^{d\times d}$ denote the unknown quantum state and let $A_{1},\dotsc,A_{m}$ be the observables in the quantum Threshold Search problem (which we assume are given in an online fashion). If we extend the unknown state $\rho$ to $\rho\otimes\lvert 0\rangle\langle 0\rvert$, then by Naimark’s Theorem 2.1, there exists a projector $\Pi_{i}\in\mathbb{C}^{d\times d}\otimes\mathbb{C}^{2\times 2}$ for each $A_{i}$ such that $\operatorname*{\mathbf{E}}_{\rho\otimes\lvert 0\rangle\langle 0\rvert}[\Pi_{i}]=\operatorname*{\mathbf{E}}_{\rho}[A_{i}]$ for all $i=1,\dotsc,m$. Since the state $\rho\otimes\lvert 0\rangle\langle 0\rvert$ can be prepared without knowing $\rho$ and this extension increases the dimension of the quantum system only by a constant factor, by replacing $\rho$ by $\rho\otimes\lvert 0\rangle\langle 0\rvert$ and each $A_{i}$ by the corresponding $\Pi_{i}$, it follows that we can assume, without loss of generality, that the observables $A_{1},\dotsc,A_{m}$ are projectors.
@@
Proof 4.3 (Proof of Theorem 4.1, assuming Lemma 4.2.).
The algorithm in Lemma 4.2 will be used as a kind of “subroutine” for the main theorem. Our first step is to augment this subroutine in the following way:
@@
We want to show that the algorithm does not make errors of type FP or FN with probability at least $0.1$. To this end, we introduce the following notation. {notation} For $i=1,\dotsc,m$, let:
@@
Therefore, since $53/4<\exp(-1/4)<4/$ 5, we have $q_{t}\leq 4/5$ and $q_{t-1}\geq 3/4$, for $\operatorname*{\mathbf{E}}[{\boldsymbol{X}}]=\Omega(\sqrt{n})$.
Assuming $\operatorname*{\mathbf{E}}[{\boldsymbol{X}}]=\Omega(\sqrt{n})$, let $t\in[m]$ be as in Claim 1. Since $q_{m}\leq q_{t}\leq 4/5$, it follows that the probability the algorithm makes an FN error is at most $4/5$. In fact, since $q_{t}\leq 4/5$, the algorithm will pick an index $i\leq t$ with probability at least $1/5$. Thus, to show that the algorithm succeeds w.p. at least $0.1$, it suffices to show that w.h.p. the algorithm does not pick an index $i\in\mathcal{B}$, where $\mathcal{B}\subseteq[m]$ is the subset defined by
@@
Proof 4.6 (Proof of Lemma 4.2).
Fix $\eta=0.01$, so that $n=O(\log^{2}m)$ as promised. By Lemma 2.10,
@@
By Claim 2 and the Cauchy–Schwarz inequality,
@@
where $i\not\in\mathcal{B}$ denotes $i\in[t]\setminus\mathcal{B}$. By Claim 1, $p_{1}+\dotsb+p_{t}\leq 1/4$. Hence,
@@
Since $\operatorname{F}(\rho^{\otimes n},\varrho_{t})\leq 1$, $\eta=0.01$, and, by Claim 1, $q_{t}\leq 4/5$, it follows that
@@
Since $\sum_{i\not\in\mathcal{B}}s_{i}$ is the probability that the algorithm returns an index $i\in[t]$ with $\operatorname*{\mathbf{E}}_{\rho}[A_{i}]\geq 1/3$, it follows that the algorithm is correct with probability at least $0.03$.
@@ 5.1 Shadow Tomography
We begin by describing how to deduce our online Shadow Tomography result, Theorem 1.4, from our online Threshold Search result, Theorem 1.1. As mentioned earlier, this deduction is known [1] to follow almost immediately from a mistake-bounded learning algorithm for quantum states due to Aaronson, Chen, Hazan, Kale, and Nayak [3], described below. We will fill in a few details that are not spelled out in [1].
@@
The above theorem is similar to, but not quite the same, as “Theorem 1” in [3]. However it is easy to check that [3]’s Section 3.3 (“Proof of Theorem 1”) applies equally well to establish Theorem 5.1 above. [^5] Briefly: the RTFL/MMW algorithm will only do an update in the “mistake” rounds. The loss is taken to be $|\widehat{\mu}_{t}-\mu^{\prime}_{t}|$. On any mistake, we have $|\widehat{\mu}_{t}-\mu_{t}|>\frac{3}{4}\epsilon$ and $|\mu^{\prime}_{t}-\mu_{t}|\leq\frac{1}{4}\epsilon$, hence the student incurs loss at least $\frac{1}{2}\epsilon$. On the other hand, answering according to the true $\mu_{t}$ would only incur loss at most $\frac{1}{4}\epsilon$. The regret calculation bounding the number of mistakes is now the same.
@@
[…]
@@
for some particular $j$. Then our algorithm can correctly output “ $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{j}]-\theta_{j}\rvert>\frac{3}{4}\epsilon$”. Furthermore, at this point the algorithm may use the holdout copies of $\rho$ to obtain an estimate $\mu^{\prime}_{j}$ of $\operatorname*{\mathbf{E}}_{\rho}[A_{j}]$ (in the naive way) that satisfies $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{j}]-\mu^{\prime}_{j}\rvert\leq\frac{1}{4}\epsilon$ except with probability at most $\delta/2$, provided $c$ is large enough.
@@
To begin, recall that our Shadow Tomography algorithm receives the input quantum events $A_{1},A_{2},\dots$ in an online fashion. As it receives them, it will run the following online algorithms concurrently:
@@
The algorithm simulates both the teacher and student roles of the mistake-bounded setting of [3] and runs in rounds. A new round is started whenever the teacher declares a mistake and a fresh batch of $n_{0}$ copies of the state $\rho$ is used by the teacher in each round. When it receives input $A_{t}$, the algorithm runs the next iteration of the mistake-bounded learning algorithm of Theorem 5.1 to get the student’s prediction $\widehat{\mu}_{t}$. Then it runs the next iteration of the Threshold Search algorithm from Lemma 5.2 with input ($A_{t}$, $\widehat{\mu}_{t}$); the estimates $\widehat{\mu}_{t}$ output by the student serve as the $\theta_{t}$ threshold values used in Lemma 5.2.
@@
It remains to verify that — assuming correct answers from all uses of Lemma 5.2 — our Shadow Tomography algorithm satisfies the Teacher Properties of Theorem 5.1 and also that all $m$ estimates for $\operatorname*{\mathbf{E}}_{\rho}[A_{i}]$ produced by the algorithm are correct to within an additive error $\epsilon$. Let us first note that within each round of the Shadow Tomography algorithm, we never supply more than $m$ quantum events to the Threshold Search algorithm from Lemma 5.2. The main point to observe is that if our Threshold Search routine from Lemma 5.2 ever passes on some $(A_{t},\widehat{\mu}_{t})$ pair, it must be that $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{t}]-\widehat{\mu}_{t}\rvert\leq\epsilon$; the reason is that passing implies the Threshold Search algorithm is prepared to output “ $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{i}]-\theta_{i}\rvert\leq\epsilon$ for all $i$”. On the other hand, it’s immediate from Lemma 5.2 that if the teacher declares “mistake” on some $(A_{t},\widehat{\mu}_{t})$ pair, then indeed we have $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{t}]-\widehat{\mu}_{t}\rvert>\frac{3}{4}\epsilon$, and the supplied correction $\mu^{\prime}_{t}$ satisfies $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{t}]-\mu^{\prime}_{t}\rvert\leq\frac{1}{4}\epsilon$ (as is necessary for the Teacher Properties, and is more than sufficient for the Shadow Tomography guarantee).